Find the minimum value of the function 

f(x, y) = x2 + 2y2 on the circle x2 + y2 = 1. 


1
Expert's answer
2021-05-07T12:16:37-0400

Solve equations "∇f= λ ∇g \text{ and } g(x,y)=1" using Lagrange multipliers


Constraint: "g(x, y)= x^2+y^2=1~~~~~~~-(11)"


Using Lagrange multipliers,


"f_x= \u03bbg_x,f_y= \u03bbg_y,g(x,y) = 1"


which become

"2x= 2xλ ~~~~~~~~~~-(1)
\\
 4y= 2yλ ~~~~~~~~~~~-(2)
\\
 x^2+y^2= 1 ~~~~~~~~~~~-(3)"


 From (1) we have "x=0 \text{ or } λ=1" . If x=0, then (11) gives y=±1. If

λ=1, then y=0 from 2, so then (11) gives x=±1. Therefore f

has possible extreme values at the points "(0,1), (0,-1), (1,0),\n\n(1,0)."


Evaluating f at these four points, we find that

• f(0,1)=2

• f(0,-1)=2

• f(1,0)=1

• f(-1,0)=1


Therefore the maximum value of f on the circle

"x^2+y^2=1 \text{ is } f(0,±1) =2" and the minimum value is "f(\u00b11,0) =1."